OpenStudy (anonymous):

write equation for slope intercept form x-intecept=-8 y-intercept=5

OpenStudy (aroub):

y-y1=m(x-x1) this is the slope intercept form right?

OpenStudy (aroub):

ok i guess it is so u just substitute the x-intercept -8 by y1 and the y-intercept 5 by x1 so the equation is y-8=m(x1+5)